Sally brought 3 chocolate bars and a pack of gum and paid $1.75. Jake brought 2 chocolate bars and 4 packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum. Thanks William
Answer by rcmcc(152)

let the chocolate bars be X
Let gum be Y
form 2 linear equations to find for x and y,
3X+Y=1.75
2X+4Y=2.00
In this case it would be easier and more simple to solve for x first,
multiply the first equation by 4, then subtract the equations to form a single linear equation
12X+4Y=7.00
2x+4Y=2.00
_____________
10X=5
divide out 10 to isolate x
X=0.50
substitute x into the first equation to solve for y,
3(.50)+Y=1.75
1.50+y=1.75
Y=.25

So a chocolate bar costs $0.50 and a pack of gum is $0.25